### v1

# LDMol: Text-Conditioned Molecule Diffusion Model Leveraging Chemically Informative Latent Space

Abstract
With the emergence of diffusion models as the frontline of generative models, many researchers have proposed molecule generation techniques using conditional diffusion models. However, due to the fundamental nature of a molecule, which carries highly entangled correlations within a small number of atoms and bonds, it becomes difficult for a model to connect raw data with the conditions when the conditions become more complex as natural language. To address this, here we present a novel latent diffusion model dubbed LDMol, which enables a natural text-conditioned molecule generation. Specifically, LDMol is composed of three building blocks: a molecule encoder that produces a chemically informative feature space, a natural language-conditioned latent diffusion model using a Diffusion Transformer (DiT), and an autoregressive decoder for molecule re. In particular, recognizing that multiple SMILES notations can represent the same molecule, we employ a contrastive learning strategy to extract the chemical informative feature space. LDMol not only beats the existing baselines on the text-to-molecule generation benchmark but is also capable of zero-shot inference with unseen scenarios. Furthermore, we show that LDMol can be applied to downstream tasks such as molecule-to-text retrieval and text-driven molecule editing, demonstrating its versatility as a diffusion model.
1 Introduction
Designing compounds with the desired characteristics is the essence of solving many chemical tasks. Inspired by the rapid development of generative models in the last decades, de novo molecule generation via deep learning models has been extensively studied. Diverse models have been proposed for generating molecules that agree with a given condition on various data modalities, including string representations [44], molecular graphs [28], and point clouds [20]. The attributes controlled by these models evolved from simple chemical properties [33, 12] to complex preferences such as biological activity [48] and multi-objective conditioning [26, 5]. More recently, as AI models’ natural language comprehension ability has rapidly increased, molecule generation aims to be controlled by natural language [29, 49].
Meanwhile, diffusion models [46, 19] have emerged as a frontline of generative models over the past few years. Through a simple and stable training objective of predicting noise from noisy data [19], diffusion models have achieved highly realistic and controllable data generation [8, 22, 18]. Furthermore, leveraging that the score function of the data distribution is learned in their training [47], state-of-the-art image diffusion models enabled various applications on the image domain [43, 23, 7]. Inspired by the success of diffusion models, several papers suggested diffusion-based molecule generative models on various molecule domains including a molecular graph [31], strings like Simplified Molecular-Input Line-Entry System (SMILES) [13], and point cloud [20].
Figure 1: Different strategies of data domain selection for molecule generative models. (a) The model directly learns the raw representation of the molecule such as Simplified Molecular-Input Line-Entry System (SMILES) tokens, point cloud coordinates, etc. (b) An autoencoder can be employed to let the generative model learn the latent distribution, but these latent spaces are generally hard to learn without any regularization on the autoencoder. (c) A regularized, chemically pre-trained encoder can provide a latent space learnable by external generative models.
However, most molecule diffusion models presented so far have used a few, relatively simple conditions to control. Also, the diffusion model was mostly trained on a raw data representation (Figure 1-(a)). In any molecule modality, the molecule’s vast and complex correlations are encapsulated between relatively small dimensions compared to common data domains like images. As a result, the output of the generative model trained on raw molecule data often failed to faithfully follow the given conditions or showed poor data quality (e.g., low validity) as the condition became more sophisticated like natural language. Furthermore, raw molecule modality has inevitable discreteness (e.g., atom and bond type, connectivity, SMILES tokens) which makes it hard to connect the deeply studied concept of continuous diffusion models to molecule generation. Naively converting discrete values into real numbers wouldn’t provide rich information within those components, and still retain certain restrictions on the data.
These issues lead us to the idea of molecule generation via latent diffusion models, which have two major advantages: Building a diffusion model on a latent space allows the model to move out of the regulated raw data domain and work on in a minimally restricted continuous domain. Furthermore, by having the encoder pre-construct a chemically useful and interpretable feature space, the diffusion model can more easily connect the molecule data with the highly complicated condition of natural texts. In fact, we show that a naive autoencoder cannot provide a feature space that the diffusion model can learn (Figure 1-(b)), and making the encoder provide a chemically informative molecule feature (Figure 1-(c)) is essential to the performance of the latent diffusion model (see Section 4.3).
Therefore, we proposed a Latent Diffusion Molecule generative model (LDMol) that enables a text-conditioned molecule generation, trained on the latent space of the separately pre-trained molecule encoder. To construct a feature space that is chemically meaningful and easily learnable, we suggest a novel contrastive learning strategy of training the encoder to encode a structural similarity of the molecule. The experimental results show that LDMol can generate molecules that are valid and also meet the input text condition. Additionally, utilizing the modeled score function, LDMol has the potential to be applied to different multi-modal downstream tasks such as molecule-to-text retrieval, text-guided molecule editing, etc. We summarize the contribution of this work as follows:
  • We propose a latent diffusion model LDMol for text-conditioned molecule generation to generate more valid molecules that are better aligned to the text condition. This approach demonstrates the potential of generative models for chemical entities in a latent space.
  • We report the importance of preparing a chemically informative latent space for the molecule latent diffusion model, and suggest a novel contrastive learning method to train an encoder that captures the molecular structural similarity.
  • LDMol outperforms the text-to-molecule generation baselines, and its modeled conditional distribution allows LDMol to retain the advanced attributes of diffusion models including various applications like molecule-to-text retrieval and text-driven molecule editing.
2 Background
Diffusion generative models. Diffusion models first define a forward process that perturbs the original data, and generates the data from the known prior distribution by the learned reverse process of the pre-defined forward process. For example, DDPM [19] fixed their forward process by gradually adding Gaussian noise to the data, which can be formalized as follows:
$q(x_{t}|x_{t-1})$ $=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$ (1)
where $\beta_{t},t=1,\dots,T$ is a noise schedule. This definition of forward process allows us to sample $x_{t}$ directly from $q(x_{t}|x_{0})$ as follows, where $\alpha_{t}=1-\beta_{t}$ and $\overline{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$:
$x_{t}$ $=\sqrt{\overline{\alpha}_{t}}x_{0}+\sqrt{1-\overline{\alpha}_{t}}\epsilon\text{, where }\epsilon\sim\mathcal{N}(0,I)$ (2)
The model $\epsilon_{\theta}$ learns the reverse process $p(x_{t-1}|x_{t})$ by approximating $q(x_{t-1}|x_{t})$ with a Gaussian distribution $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\sigma_{t}^{2}I)$ where
$\mu_{\theta}(x_{t},t)$ $=\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\overline{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t)\right)$ (3)
which can be trained by minimizing the difference between $\epsilon$ and $\epsilon_{\theta}(x_{t},t)$:
$\theta^{*}$ $=\text{arg}\min_{\theta}\mathbb{E}_{x_{0},t,\epsilon}||\epsilon-\epsilon_{\theta}(x_{t},t)||_{2}^{2}$ (4)
Once $\theta$ is trained, novel data can be generated with the learned reverse process $p_{\theta}$: starting from the random noise $x_{T}\sim\mathcal{N}(0,{I})$, the output can be gradually denoised according to the modeled distribution of $p_{\theta}(x_{t-1}|x_{t})$.
Various real-world data generation tasks require to generate data $x_{0}$ with a given condition $c$. To build diffusion models that can generate data from the conditional data distribution $q(x_{0}|c)$, the model that predicts the injected noise should be also conditioned by $c$.
$\theta^{*}$ $=\text{arg}\min_{\theta}\mathbb{E}_{x_{0},y,t,\epsilon}||\epsilon-\epsilon_{\theta}(x_{t},t,c)||_{2}^{2}$ (5)
Although the reverse diffusion process with ancestral sampling of $x_{0}\sim q(x_{0}|y)$ can be identical to unconditional diffusion models,  [18] proposed classifier-free guidance to improve the quality of the conditional diffusion models’ output. Specifically, using a scalar guidance scale $\omega$, classifier-free guidance sampling utilizes a linearly combined noise of conditional and unconditional model output,
$\hat{\epsilon_{\theta}}(x_{t},t,c)=\epsilon_{\theta}(x_{t},t,\emptyset)+\omega\{\epsilon_{\theta}(x_{t},t,c)-\epsilon_{\theta}(x_{t},t,\emptyset)\}$ (6)
where $\emptyset$ denotes the special condition input for unconditional noise.
As the diffusion model’s target data domain becomes more complex and high-dimensional, diffusion models began to perform their diffusion process on the latent data space instead of the raw data [51, 39]. Specifically, using a pre-trained pair of an encoder $\mathcal{E}$ and its corresponding decoder $\mathcal{D}$, datapoint $x$ is encoded into $z:=\mathcal{E}(x)$ that is used to train the diffusion model. Generated latent $z^{\prime}$ from the pre-trained diffusion model is transformed back to $x^{\prime}:=\mathcal{D}(z^{\prime})$ on the original data domain.
Conditional molecule generation. Compared to more common domains for generative models like images or text, generative models for molecules are more challenging since there are rigid constraints on the data modality itself. String representations like SMILES has a strict syntax, which often makes SMILES-based molecule generative models show low validity on their output. A similar challenge also exists for the modalities of point clouds and graphs to ensure the model output represents a valid molecule.
As a promising tool for many important chemical and engineering tasks like de novo drug discovery and material design, conditional molecule generation has been extensively studied with various models including recurrent neural networks [44], graph neural networks [28], and variational autoencoders [14, 27]. With the advent of large and scalable pre-trained models with transformers [53], the controllable conditions became more abundant and complicated [1, 5], and reached a text-guided molecule generation [10, 49, 29] which requires a deep comprehension ability for natural language.
Recent works attempted to import the success of the diffusion model into molecule generation. Several graph-based and point cloud-based works have built conditional diffusion models that could generate molecules with simple physicochemical and biological properties [20, 31, 50]. The most related model to our work would be TGM-DLM [13], which attempted a text-conditioned molecule diffusion model trained on the sequence of tokenized SMILES indices.
3 Method
3.1 Chemically informative SMILES authoencoder
The primary goal of introducing autoencoders for image latent diffusion models is to map raw images into a low-dimensional space which reduces the computation cost [51, 39]. This is plausible because a high-resolution image has an enormous dimension on a pixel domain, yet each pixel contains little information.
In molecule generation, we chose the SMILES string as a base modality to represent molecules, and built a SMILES encoder to map raw SMILES strings into a latent vector. In this case, the role of our SMILES autoencoder has to be different from the autoencoders for images. A molecule structure can be fully expressed by a sequence of $L$ integers for SMILES tokens, where $L$ is the maximum token length. However, each token carries a significant amount of information, and hidden interactions between these tokens are much more complicated than interactions between image pixels. Therefore, the SMILES encoder should be more concerned with extracting chemical meaning into the latent space rather than reducing dimensions, even if it results in a latent space whose dimension is larger than the raw SMILES string. In fact, when a SMILES autoencoder was trained with naive reconstruction loss only, we observed that diffusion models failed to learn the latent space from this autoencoder (as will be seen in Section 4.3). This phenomenon also occurs even in the image latent to some extent, which is why autoencoders for image diffusion models are usually trained with regularizations such as KL divergence loss [17] or vector quantization [52].
Accordingly, here we propose three conditions that our SMILES autoencoder’s latent space has to satisfy: the input should be reconstructible, embed chemical information that diffusion models can learn, and have a small dimension as possible. A number of molecule encoders have been presented that can extract various useful chemical features [54, 29] including biochemical activity or human-annotated descriptions [55, 30]. Nonetheless, these molecule feature extractors didn’t aim to encode the entire information of the molecule structure therefore the input cannot be restored from the model output. Considering that an input molecule has to be reconstructed from the latent, the most important characteristic that our SMILES latent should have is that all SMILES representing the same molecule must be encoded into the same vector.
Figure 2: Overview of the proposed molecule autoencoder and the latent diffusion model. (a) SMILES encoder is trained with contrastive learning to construct latent space that embeds a structural similarity. (b) After the SMILES encoder is prepared, a compression linear layer and the autoregressive decoder are trained to restore the encoder input. (c) Training and inference process of the latent diffusion model conditioned by the output of the frozen external text encoder.
Encoder design. In this respect, we trained our SMILES encoder with contrastive learning similar to simCLR [6](Figure 2-(a)). Contrastive learning aims to learn better representation by assimilating features containing similar information (i.e. positive pair) and distancing semantically unrelated features (i.e. negative pair). Taking advantage of the SMILES’s characteristic that a single molecule can be represented with many different SMILES strings, we defined two different SMILES from the same molecule as a positive pair. In the process of matching the outputs of two SMILES in a positive pair, the encoder learns the syntax and the information within SMILES so that its output encapsulates the structural information of the input molecule.
Specifically, a SMILES string is fed into the SMILES encoder with a special "[SOS]" token which denotes the start of the sequence. For a batch of $N$ input SMILES $M=\{m_{1},m_{2},\dots,m_{k},\dots,m_{N}\}$, we prepare a positive pair SMILES $m_{k}^{\prime}$ for each $m_{k}$ by SMILES augmentation to construct $M^{\prime}=\{m_{1}^{\prime},m_{2}^{\prime},\dots,m_{k}^{\prime},\dots,m_{N}^{\prime}\}$. After $M$ and $M^{\prime}$ are passed through the SMILES encoder, we fed each SMILES’ output vector corresponding to the [SOS] token into an additional linear projection and normalization layer, denoting the output features as $v_{k}$ and $v_{k}^{\prime}(k=1,2,\dots,N)$.
Assimilating $v_{k}$ and $v_{k}^{\prime}$ from the positive pairs and distancing the others can be done by minimizing the following contrastive loss, which is essentially cross-entropy loss where the logit is the cosine similarity between given pairs.
$\mathcal{L}_{contrastive}(M,M^{\prime})=-\sum_{k=1}^{N}\log\frac{\exp(v_{k}\cdot v_{k}^{\prime}/\tau)}{\sum_{i=1}^{N}\exp(v_{k}\cdot v_{i}^{\prime}/\tau)}\;$ (7)
Here, $\tau$ is a positive temperature parameter. To utilize a symmetric loss against the input, we trained our encoder with the following loss function.
$\mathcal{L}_{enc}(M,M^{\prime})=\mathcal{L}_{contrastive}(M,M^{\prime})+\mathcal{L}_{contrastive}(M^{\prime},M)\;$ (8)
While our encoder was pre-trained to map structurally similar molecules to similar vectors, we found that the encoder tried to map stereoisomers to almost the same point in the latent space; indeed, in terms of structural differences, stereoisomers are the most challenging molecules to distinguish. To resolve this, if the training data has a stereoisomer, we considered it as ”hard-negative" samples and included it in the loss calculation batch with the original data.
Compressing the latent space. The pre-trained SMILES encoder maps a molecule into a vector of size $\{L\times d_{enc}\}$, where $d_{enc}$ is a feature size for the encoder. To avoid the curse of dimensionality and construct a more learnable feature space for diffusion models, we additionally employed a linear compression layer (Figure 2-(b)) to reduce the dimension size from $\{L\times d_{enc}\}$ to $\{L\times d_{z}\}$. The range of this linear layer output $z$ is a target domain of our latent diffusion model.
Decoder design. We prepared the SMILES decoder to reconstruct the input SMILES sequence from a given $z$. Following many major works that treated SMILES as a variant of language data[44, 5, 41], we built an autoregressive transformer [53] as our decoder (Figure 2-(b)) which is widely used to successfully generate sequential data [4]. Specifically, starting from the [SOS] token, the decoder predicts the next SMILES token using information from $z$ with cross-attention layers. When $m=\{t_{0},t_{1},\dots,t_{n}\}$ is the token sequence of the input SMILES, the decoder is trained to minimize the next-token prediction loss described as Eq. (9):
$L_{dec}=-\sum^{n}_{i=1}\log p(t_{n}|t_{0:n-1})\;$ (9)
Here, the decoder and the compression layer are jointly trained while the encoder’s parameter is frozen. After being fully trained, the decoder was able to reconstruct roughly 98% of the SMILES encoder input.
3.2 Text-conditioned latent diffusion model
As shown in Figure 2-(c), our diffusion model learns the conditional distribution of the SMILES latent $z$ whose dimension size is $\{L\times d_{z}\}$. In the training phase, a molecule in the training data is mapped to the latent $z$ and applied a forward noising process into $z_{t}$ with randomly sampled diffusion timestep $t$ and injected noise $\epsilon$. A diffusion model predicts the injected noise from $z_{t}$, conditioned by the paired text description via a frozen external text encoder. In the inference phase, the diffusion model iteratively generates a new latent sample $z^{\prime}$ from a given text condition, which is then decoded to a molecule via the SMILES decoder.
Since most contributions to diffusion models were made in the image domain, most off-the-shell diffusion models have the architecture of convolution-based Unet [40]. However, to step beyond the spatial inductive bias and make the model applicable for more general data modality, diffusion models using transformer architectures are proposed. One of the most successful approaches among them is DiT [35], which is the model we’ve employed. Specifically, we utilized a DiTbase model with two major modifications: first, DiT’s two-dimensional attributes (e.g., patch embedding) were modified accordingly to handle one-dimensional latent. Second, cross-attention layers for external conditions were added after the self-attention layers to receive high-dimensional text conditions.
Text-based molecule generation requires a text encoder to process natural language conditions. Existing text-based molecule generation models trained their text encoder from scratch [36], or utilize a separate encoder model pre-trained on scientific domain corpora [3]. In this work, we took the encoder part of MolT5 [10] as our text encoder.
3.3 Implementation details
The pre-training of the SMILES encoder and the corresponding decoder was done with 10,000,000 general molecules from PubChem [24]. To ensure enough batch size for negative samples [15], we build a memory queue that stores $Q$ recent input and use them for the encoder training. To be fed into the autoencoder, an input SMILES must be tokenized as a sequence of tokens in a pre-defined vocabulary. The tokenizer vocabulary consists of 300 tokens, which were obtained from the pre-training data SMILES corpus using the BPE algorithm [11]. We only used SMILES data that does not exceed a fixed maximum token length $L$.
To train the text-conditioned latent diffusion model, we gathered three existing datasets of text-molecule pair: PubchemSTM curated by [30], ChBEI-20 [9], and PCdes [55]. Only a train split for each dataset was used for the training, and pairs that appear in the test set for the experiments in the Results are additionally removed. We also used 100,000 molecules from PubChem without any text descriptions, which helps the model learn the common distribution of molecules. When these unlabeled data were fed into the training model, we used a pre-defined null text which represents the absence of text condition.
The latent diffusion model was trained with the training loss suggested in ADM [8]. To take advantage of classifier-free guidance, we randomly replaced 3% of the given text condition with a null text during the training. More detailed training hyperparameters can be found in Appendix A. Unless specifically mentioned, the sampling iteration in the inference stage used DDIM-based [45] 100 sampling steps with a classifier-free guidance scale of 5.0.
4 Experiments
4.1 Text-conditioned molecule generation
In this section, we evaluated the trained LDMol’s ability to generate molecules that agree with the given natural language conditions. We designed two scenarios: a benchmark study to see if the model can find the output that satisfies very long and specific conditions, and zero-shot text-to-molecule generation with unseen text conditions to see LDMol’s generalization ability and the learned alignment between text and molecules.
Quantitative benchmark study. We generated molecules with LDMol using the text captions in the ChEBI-20 test set and compared them with the ground truth. Here, we assume that the given description is long and detailed so that the ideal generative model can figure out the ground truth. On top of the training with molecule-text pairs from various sources, we additionally fine-tuned the model with the ChEBI-20 training set and 10,000 unlabeled molecules. The metrics we’ve used are as follows: SMILES validity, BLEU score [34] and Levenshtein distance between two SMILES, Tanimoto similarity [2] between two SMILES with three different fingerprints (MACCS, RDK, Morgan), the exact match ratio, and Frechet ChemNet Distance (FCD) [37].
We note that since our diffusion model sampling starts from randomly sampled noise from a prior distribution, the output from deterministic DDIM sampling still has stochasticity. However, the performance difference was marginal between starting from random noise and starting from a fixed point (e.g., zero vector). Since the given condition is more specific and detailed than zero-shot case studies, we used the lower classifier-free guidance scale of 3.5.
Table 1: Benchmark results of text-to-molecule generation on ChEBI-20 test set. The best performance for each metric was written in bold, and the second-best performance was underlined. The results for baseline models are taken from  [13].
[…]
Transformer 0.906 0.499 57.660 0.480 0.320 0.217 0.000 11.32
SciBERT [3] 0.915 0.459 24.458 0.721 0.588 0.529 0.005 -
T5base [38] 0.660 0.765 24.950 0.731 0.605 0.545 0.069 2.48
MolT5base [10] 0.772 0.769 24.458 0.721 0.588 0.529 0.081 2.18
GIT-Mol [29] 0.928 0.756 26.315 0.738 0.582 0.519 0.051 -
TGM-DLM [13] 0.871 0.826 17.003 0.854 0.739 0.688 0.242 0.77
LDMol 0.928 0.900 9.259 0.931 0.860 0.820 0.363 0.25
Table 1 contains the performance of LDMol and other baselines for text-to-molecule generation on the ChEBI-20 test set. Including both transformer-based autoregressive models and diffusion-based models, LDMol outperformed the existing models in almost every metric. While one model showed similar validity to ours, it had a poor agreement with the ground truth compared to LDMol, with an exact match of 0.051. One thing to emphasize is that MolT5 uses the same text encoder with LDMol, yet there’s a significant performance difference between the two models. We believe this is because our constructed latent space is much easier to align with the same textual information, compared to the token sequence used by transformer-based models.
Figure 3: Examples of the generated molecules by LDMol with various text conditions. Validity on 1,000 generated samples is included in the lower left corner of each box.
Zero-shot case studies. To demonstrate the generalization ability of LDMol, we analyzed the generated output with several hand-written text conditions. These input prompts were not contained in the training data and were relatively vague and high-level so that many different molecules could satisfy the condition.
Figure 3 shows the samples of generated molecules from LDMol with several input prompt examples. We found that LMDol can generate molecules with high validity that follow the various levels of input conditions for specific atoms(a), functional groups(b), molecular substructure(c), compound class(d), and substance names(e). The validity was calculated by the number of valid SMILES over the 1,000 generated samples, which was above 0.9 for all five scenarios we’ve shown. Considering that these short, broad, and hand-written text conditions are distinct from the text conditions in the training dataset, we’ve concluded that our model is able to learn the general relation between natural language and molecules.
4.2 Applications toward downstream tasks
Well-trained diffusion models learned the score function of a data distribution, which enables high applicability to various downstream tasks. The state-of-the-art image diffusion models have shown their versatility in image editing [32, 16], classification [25], retrieval [21], inverse problems like inpainting and deblurring [7], image personalization [42], etc. To demonstrate LMDol’s potential versatility as a diffusion model, we applied the pre-trained LDMol to the molecule-to-text retrieval and text-driven molecule editing task.
Molecule-to-text retrieval. Our approach to molecule-to-text retrieval is similar to the idea of using a pre-trained diffusion model as a classifier [25]; LDMol takes each candidate text with a query molecule’s noised latent, and retrieves the text that minimizes the difference between output noise $\hat{\epsilon_{\theta}}$ and the injected noise $\epsilon$ (see Appendix B for more details). Since this process has randomness due to the stochasticity of $t$ and $\epsilon$, we repeated the same process $n$ times with resampled $t$ and $\epsilon$ and used a mean loss to minimize the performance variance.
Table 2: 64-way accuracy of molecule-to-text retrieval task. The custom test set and the baseline model results are taken from  [49]. For LDMol, $n$ is a number of iterations where $||\hat{\epsilon_{\theta}}-\epsilon||^{2}_{2}$ was calculated. The best performance for each task is written in bold.
Model PCdes test set custom test set from Su et al.
sentence-level paragraph-level sentence-level paragraph-level
SciBERT [3] 50.38 82.57 1.38 1.38
KV-PLM [55] 55.92 77.92 1.37 1.51
MoMu-S [49] 58.64 80.62 39.50 45.66
MoMu-K [49] 58.74 81.09 39.07 46.21
LDMol($n$=10) 52.28 89.36 57.90 78.16
LDMol($n$=25) 55.94 90.82 63.27 81.13
We measured a 64-way retrieval accuracy of LDMol using two different test sets: PCdes test split and another text-molecule retrieval dataset separately curated by  [49]. Table 2 contains the text-to-molecule accuracy of LDMol with other baseline models. Only one randomly selected sentence in each candidate description was used for the retrieval in the “sentence-level" column, and the entire descriptions were used for the “paragraph-level" column. LDMol achieved a higher performance in three out of four scenarios compared to the previously presented models and became more accurate as the number of function evaluations increased. It maintained its performance on a relatively out-of-distribution test set with minimal accuracy drop. The actual examples from the retrieval result can be found in Appendix C.
Text-driven molecule editing. We applied a method of Delta Denoising Score (DDS) [16], which was originally suggested for text-guided image editing, to see whether LDMol can be used to optimize a source molecule to match a given text. Using two texts that describe the source data $z_{src}$ and the desired target, DDS presents how a text-conditioned diffusion model can modify $z_{src}$ to a new data $z_{tgt}$ that follows the target text description (more detailed procedures can be found in Appendix B).
Figure 4: Examples of text-guided image editing with LDMol. The difference between the source text and the target text, and the corresponding region is colored in purple.
We imported a method of DDS on LDMol’s molecule latent to edit a given molecule to match the target text, where the case studies are shown in Figure 4., The editing results successfully modified the input molecule towards the target text with minimal corruption of the unrelated region. We emphasize that this kind of direct data optimization is possible because of the continuous and non-restricted latent domain from the LDMol’s encoder.
4.3 Ablation studies
We’ve conducted ablation studies for each design choice of the proposed LDMol to analyze and emphasize their role. We measured the SMILES validity and the exact match ratio of text-to-molecule generation on the ChEBI-20 test set, with three models pre-trained with the same number of DiT training iterations.
Table 3: Quantitative results of the ablation study.
[…]
LDMol 0.922 0.303
w/o contrastive encoder pre-training $0$
w/o compression layer 0.110 0.003
The result of the ablation study is listed in Table 3. When the molecule latent space was constructed by a naive autoencoder (i.e. the SMILES encoder is not pre-trained with the proposed method), the diffusion model completely failed to learn the data distribution to generate valid SMILES. A similar problem occurred when we didn’t introduce a compression layer to reduce the dimension of the latent space, meaning that the diffusion model could not properly model the data distribution in the given latent space.
5 Conclusion
In this work, we presented a text-conditioned molecule diffusion model LDMol that runs on a chemical latent space reflecting structural information. By introducing the deeply studied paradigm of the latent diffusion model in general domains into the molecule domain with minimal modifications, LDMol retains many advanced attributes of diffusion models that enable various applications.
Despite the noticeable performances of LDMol, it still has limitations that can be improved; Since the text captions of our obtained paired data mostly describe structural features, LDMol still struggles to follow some text conditions such as biological properties. The current suggested text-guided molecule editing performance is relatively sensitive to hyperparameters as the imported method assumes broader and more substantial knowledge that emerged from the data generation. Nonetheless, we expect that the LDMol’s performance could be improved further with the emergence of richer text-molecule pair data and more powerful text encoders, with possible future works for the construction of a better latent space.
We believe that our approach could inspire tackling various chemical generation tasks using latent space, not only text-conditioned but also many more desired properties such as biochemical activity. Especially, we expect LDMol could be a starting point to bridge achievements in the state-of-the-art diffusion model into the chemical domain.
Appendix A Model hyperparameters and training setup
Table 4: Our choice of the model hyperparameters and training setup.
hyperparameters
$L$ 128
$d_{enc}$ 1024
$d_{z}$ 64
$\tau$ 0.07
$Q$ 16384
training setup
optimizer autoencoder: AdamW, DiT: Adam
learning rate autoencoder: cosine annealing(1e-4$\rightarrow$1e-5), DiT: 1e-4
batch size per gpu encoder: 64, decoder: 128, DiT: 16
training resources encoder: NVIDIA A100(VRAM:40GB), 15 hours decoder: NVIDIA A100(VRAM:40GB), 15 hours DiT: 8 NVIDIA A100(VRAM:40GB), 96 hours
Appendix B More detailed overview for LDMol’s application for downstream tasks
Figure 5: Overall pipeline for the applications of LDMol. (a) Molecule-to-text retrieval. (b) Text-guided molecule editing. The SMILES autoencoder and the text encoder are not drawn in this figure.
Figure 5-(a) shows the LDMol’s molecule-to-text retrieval process. A given query molecule is converted to a latent $z$ then a forward noise process is applied with a randomly sampled timestep $t$ and noise $\epsilon$. This $z_{t}$ is fed to LDMol with each candidate text, and the candidate that minimizes the loss $||\hat{\epsilon_{\theta}}-\epsilon||^{2}_{2}$ between $\epsilon$ and the output noise $\hat{\epsilon_{\theta}}$ is retrieved. To minimize the variance from the stochasticity of $t$ and $\epsilon$, the same process can be repeated multiple times with resampled $t$ and $\epsilon$ to use a mean loss.
Figure 5-(b) illustrates the DDS-based molecule editing with LDMol. Specifically, DDS requires source data $z_{src}$, target data $z_{tgt}$ which is initialized to $z_{src}$, and their corresponding source and target text descriptions. We apply the forward noise process to $z_{src}$ and $z_{tgt}$ using the same randomly sampled $t$ and $\epsilon$ to get $z_{t,src}$ and $z_{t,tgt}$. These are fed into the pre-trained LDMol with their corresponding text, where we denote the output noise as $\hat{\epsilon}_{\theta,src}$ and $\hat{\epsilon}_{\theta,tgt}$. Finally, $z_{tgt}$ is modified towards the target text by optimizing it to the direction of $(\hat{\epsilon}_{\theta,tgt}-\hat{\epsilon}_{\theta,src})$ with a learning rate $\gamma$. Here, $\hat{\epsilon}_{\theta,tgt}$ and $\hat{\epsilon}_{\theta,tgt}$ can be replaced with the classifier-free-guidanced noises, using the output with the null text and the guidance scale $\omega$. We repeated this optimization loop 200 times with $\gamma=0.4$ and $\omega=2.0$.
Appendix C Molecule-to-text retrieval examples
Figure 6: The examples of molecule-to-text retrieval result on the PCdes test set. The retrieval was done at the sentence level, and the top three retrieval outputs for each query molecule are described. The corresponding description from the data pair was correctly retrieved at first for both cases, and the other retrieved candidates weakly show a correlation with the query molecule.